Stan and Hilda can mow the lawn in 40 min if they work together. If Hilda works twice as fast as Stan, how long would it take Stan to mow the lawn alone?
Answer by stanbon(75887)

Stan and Hilda can mow the lawn in 40 min if they work together. If Hilda works twice as fast as Stan, how long would it take Stan to mow the lawn alone?
------------
Together rate = 1/40 job/min
Hilda time = x: Hilda rate = 1/x job/min
Stan time = 2x; Stan rate = 1/(2x) job/min
------------------------------
Equation:
rate + rate = together rate
1/x + 1/2x = 1/40
-----
Multiply thru by 40x to get:
40 +20 = x
x = 60 minutes
----
Stan time = 2x = 2 hrs
============================
Cheers,
Stan H.